# PersonaCraft: Personalized Full-Body Image Synthesis for Multiple Identities from Single References Using 3D-Model-Conditioned Diffusion

### v1

@@ 3.1 Face and Body Identity Extraction
$\vspace{-3pt}\mathbb{E}_{\bm{x}_{0},t,\bm{c},\bm{d}_{\text{SMPLx}},\epsilon}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{x}_{t},{t},\bm{c},\mathcal{E}^{\text{SC}}_{\phi}({\bm{d}_{\text{SMPLx}}}))\|_{2}^{2}\Big{]},$ (1)
where $\epsilon_{\theta}$ is a pretrained diffusion model, kept frozen during SCNet training, and $\epsilon\sim\mathcal{N}(0,1)$.
Figure 5: Additional comparison of generated images. Yellow arrows highlight anatomical issues due to 2D pose limitations. Blue arrows refer to the individuals evaluated for correct body shape preservation. PersonaCraft excels in identity, body shape consistency, and naturalness while being over twice as fast as OMG-based methods [71, 80, 43].
3.3 Multiple Human Personalized Image Synthesis
To synthesize personalized images of multiple individuals, we leverage a face ControlNet, IdentitiyNet  [72], and SCNet. To ensure precise identity preservation and enhanced image quality, we employ face masks to accurately localize facial regions from the given pose. Suppose we aim to generate a personalized image, given a set of $N_{\text{human}}$ reference images $\{I_{\text{ref}}^{(i)}\}^{N_{\text{human}}}_{i=1}$, corresponding target poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, and a text prompt $y$. Then, we obtain face embeddings $\{\bm{f}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and body shape parameters $\{\bm{\beta}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and corresponding SMPLx depth $\bm{d}_{\text{SMPLx}}$ following the processes in Sec. 3.1 and Sec. 3.2.
We then obtain residual features $\{R^{(i)}_{\text{face}}\}^{N_{\text{human}}}_{i=1}$ generated from the shared IdentityNet [71], $\mathcal{E}^{\text{ID}}_{\omega}$, given face embeddings $\{\bm{f}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and face landmarks $\{\bm{p}_{\text{face}}^{(i)}\}^{N_{\text{human}}}_{i=1}$. Landmarks can be extracted from reference images or provided as additional 3D pose inputs. Also the residual feature $R_{\text{body}}$ is generated with SCNet $\mathcal{E}^{\text{SC}}_{\phi}$, given text $y$ and $\bm{d}_{\text{SMPLx}}$ derived from ${\{I_{\text{ref}}^{(i)}\}}^{N_{\text{human}}}_{i=1}$ and $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$.
$\vspace{-3pt}R^{(i)}_{\text{face}}=\mathcal{E}^{\text{ID}}_{\omega}(\bm{f}^{(i)},\bm{p}_{\text{face}}^{(i)}),\ \ R_{\text{body}}=\mathcal{E}^{\text{SC}}_{\phi}(y,\bm{d}_{\text{SMPLx}})$ (2)
@@ 3.3 Multiple Human Personalized Image Synthesis
Let $F^{(k)}_{\theta}(\cdot)$ represent the $k$-th neural block and $\bm{s}^{(k)}$ the $k$-th input feature map. We obtain $\bm{s}^{(k+1)}$ by adding these residual features to $F^{(k)}_{\theta}(\bm{s}^{(k)})$, scaled by their respective conditioning weights $\alpha_{\text{face}}$ and $\alpha_{\text{body}}$, and modulated by face masks $M^{(i)}$:
@@
$\vspace{-3pt}\bm{s}^{(k+1)}=F^{(k)}_{\theta}(\bm{s}^{(k)})+\alpha_{\text{face}}\sum\limits_{i=1}^{N_{\text{human}}}M^{(i)}\ast R^{(i)}_{\text{face}}+\alpha_{\text{body}}R_{\text{body}}$ (3)
Unlike two-stage methods (e.g., OMG [43]) or test-time optimization approaches (e.g., DreamBooth [60]), our method is more efficient and preserves the versatility of foundation models like SDXL. This enables training-free multi-human personalization from single reference images, with scalability to accommodate numerous identities without performance degradation.
@@ 3.4 User-Defined Body Shape Control
The SMPLx [57] shape parameters $\bm{\beta}$ are multiplied elementwise with the orthonormal shape displacement components, $\mathbf{S}$, and then added to the base shape, $\bm{v}_{\text{template}}$. Given $\bm{\beta}\in\mathbb{R}^{10}$, $\mathbf{S}\in\mathbb{R}^{N_{\text{vertices}}\times 3\times 10}$, and $\bm{v}_{\text{template}}\in\mathbb{R}^{N_{\text{vertices}}\times 3}$, the shaped vertices are defined as.:
$\vspace{-3pt}\bm{v}_{\text{shaped}}=\bm{v}_{\text{template}}+\sum_{i=1}^{10}\beta_{i}\cdot\mathbf{S}[:,:,i].$ (4)
Figure 6: Additional qualitative examples of PersonaCraft for single and multi-human personalization.
Table 1: Quantitative evaluation of PersonaCraft in face identity and body shape preservation, showing superior performance over baselines. CLIP similarity is comparable, with the result of text correspondence in user study (Tab. S1) addressing CLIP’s limitations in complex scenes.
Face ID preservation↑ Body shape preservation↑ CLIP similarity↑ # of idenities 1 2 3 4 5 Total 1 2 3 4 5 Total InstantID + OMG 0.369 0.274 0.200 0.169 0.155 0.194 0.505 0.406 0.349 0.307 0.302 0.370 0.267 IP-Adapter + OMG 0.227 0.127 0.134 0.118 0.134 0.125 0.538 0.442 0.377 0.340 0.335 0.403 0.266 IPA-Face + OMG 0.310 0.222 0.160 0.145 0.136 0.160 0.479 0.404 0.345 0.308 0.312 0.367 0.267 Ours 0.404 0.357 0.317 0.277 0.267 0.305 0.592 0.575 0.510 0.502 0.481 0.531 0.252
Table 2: Top-1 preference rates from user study on naturalness, face identity, body shape preservation, and text-image correspondence.
Top 1 (%) Natural.↑ Face ID↑ Body shape↑ Text corr.↑ DreamBooth 13.01 11.38 11.15 15.21 Textual Inversion 9.41 9.64 9.99 13.36 InstantID + OMG 5.92 9.99 8.83 8.36 IP-Adapter + OMG 15.45 13.24 14.29 12.66 IPA-Face + OMG 10.57 12.78 11.85 9.18 Ours 45.64 42.97 43.90 41.23
PersonaCraft offers user-defined body shape control, allowing adjustments based on user preferences. Specifically, we provide two types of controls: 1) Reference-based body shape control: This approach derives the body shape parameter $\bm{\beta}_{\text{target}}$ from the reference image using SMPLx fitting and directly applies it. 2) Interpolation and extrapolation-based body shape control: Here, we acquire two reference body shape parameters, $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$, and generate a controlled shape by interpolating or extrapolating between them. This target shape is defined as $\bm{\beta}_{\text{target}}=(1-\gamma)\bm{\beta}_{1}+\gamma\bm{\beta}_{2}$, where $\gamma\in\mathbb{R}$ adjusts the blend between $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$. The resulting $\bm{\beta}_{\text{target}}$ is then applied to Eq. 4 to obtain the target body shape. Results are shown in Fig. 7.
4 Experiments
4.1 Experimental Details
Implementation details. We implemented PersonaCraft using Stable Diffusion XL (SDXL) [59] as the base model. SCNet was trained on a curated dataset with a learning rate of 1e-5, using the Adam optimizer and a batch size of 2. The training process ran for 50,000 iterations on a single NVIDIA A100 GPU. For all experiments, we utilized 50 denoising steps.
Baselines. To evaluate the performance of PersonaCraft, we compared it against several baseline methods for single-shot, multi-identity, and pose-controllable image synthesis. All baselines were implemented using SDXL [59] to ensure a fair comparison. Key baselines include OMG [43], a state-of-the-art multi-concept personalization method, combined with InstantID/IP-Adapter/IP-Adapter-Face (IPA-Face) [71, 80], a leading single-shot personalization model and 2D pose ControlNet [81]. Optimization-based methods such as DreamBooth [60] and Texture Inversion [23] were also included for comparison. Additional results with other baselines, which may not align perfectly with the above categories, are also provided in the supplementary.
Test dataset. For testing, we evaluated our method using the COCO-WholeBody dataset [37], which contains full-body images of multiple people. The dataset includes images of 1 to 5 people, with a total of 1,000 images selected (200 images for each number of people). Text prompts were extracted from these images using BLIP [47] to generate corresponding textual descriptions. For pose conditioning, MultiHMR [12] was used to obtain 2D poses and 3D poses.
Metrics. Face identity preservation was measured for 1-5 identities following FastComposer [75], using FaceNet [63] for identity similarity within the face mask. Body shape preservation was evaluated using cosine similarity between SMPLx $\bm{\beta}$ parameters. Text-image correspondence was measured with CLIP-L/14 for image-text alignment. Additionally, a user study was conducted to assess perceptual quality.
More experimental details are in the supplementary.
Figure 8: Results of combining PersonaCraft with various style LoRAs, showcasing adaptability to styles like Pastel, JoJo, and Pokémon Trainer. Some styles alter facial and body identities due to their bias, while producing visually impressive outcomes.
4.2 Results
Qualitative results. As shown in Fig. 45, and 6, our proposed method, PersonaCraft, demonstrates significant advantages over existing approaches, particularly in handling multi-identity images with occlusions and challenging poses. By leveraging a 3D-human model-conditioned ControlNet, PersonaCraft accurately preserves body shapes, enhances pose fidelity, and maintains body feature consistency with fewer artifacts in complex scenarios.
In contrast, methods like InstantID+OMG, IP-Adapter+OMG, and IPA-Face+OMG [71, 80, 43], which rely on 2D skeleton-based pose conditioning, exhibit severe anatomical inaccuracies in occluded or intricate interactions (highlighted by yellow arrows). These limitations arise from the inherent weaknesses of 2D pose representations, which struggle with overlapping body parts and complex scenarios. Additionally, blue arrows highlight cases where PersonaCraft maintains accurate body structure and pose fidelity, outperforming these methods in both identity preservation and body shape consistency.
DreamBooth [60] and Textual Inversion [23], which lack explicit pose guidance, often suffer from severe anatomical distortions, omission of individuals in multi-person scenes, and issue with clothing-body shape displacement, where clothing styles fail to adapt to individual body shapes.
Unlike 2D pose ControlNet, which struggles to distinguish occluded regions and handle challenging poses, PersonaCraft addresses these limitations effectively. It demonstrates robust performance as the number of subjects increases, preserving identity and body structure in multi-subject scenes. These findings highlight PersonaCraft’s robustness and versatility, establishing it as a state-of-the-art solution for personalized image generation in complex, real-world scenarios.
Quantitative results. We evaluate PersonaCraft’s performance in terms of face identity preservation, body shape preservation, and text-image correspondence. As stated in Tab. 1, the results show that our method consistently outperforms baselines in both face identity and body shape preservation across varying numbers of identities. Notably, it demonstrates robust body shape preservation with minimal variation across different body shapes. While text-image correspondence scores are comparable, it is important to note, as highlighted in previous literature [45, 42], that CLIP scores are less effective in evaluating images with complex scenes. To supplement this, we also conducted user studies, where PersonaCraft’s performance was highly rated.
User study. We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked their top three preferences for each attribute, yielding 10,332 responses from 123 participants. Results in Tab. S1 indicate that our method achieved the highest Top-1 preference across all metrics, with a significant margin over the baselines. Notably, the global CLIP score appeared misaligned with human judgment, underscoring its limitations in complex scenes. Further details of the user study setup are provided in the supplementary material.
Figure 9: Effectiveness of (a) SCNet compared to 2D Pose ControlNet [81] and (b) masking in multi-human image synthesis.
4.3 User-Defined Body Shape Control
@@
A key feature of PersonaCraft is its user-defined body shape control, as demonstrated in Fig. 7. Users can select a reference image with the desired body shape, allowing the model to adapt these characteristics to generate highly personalized images. The system offers additional flexibility by enabling users to adjust body proportions through smooth interpolation or extrapolation between two different body references. This ability to fine-tune body shapes provides enhanced customization, making PersonaCraft a powerful and versatile tool for creating tailored image outputs with precise body shape control.
Table 3: Inference times for multi-identity personalized synthesis.
Method Total Time (secs) Text Inversion 1636.15 DreamBooth 770.71 InstantID + OMG 46.94 IP-Adapter + OMG 44.62 IPA-Face + OMG 35.46 PersonaCraft (ours) 17.25
Table 4: Evaluation of body shape preservation with and without the use of the body shape parameter.
Single Multi Total w/o body shape 0.559 0.493 0.506 w/ body shape 0.592 0.517 0.531
4.4 PersonaCraft with Stylization
The proposed method is a plug-and-play approach, making it compatible with various style-specific LoRAs. To evaluate its effectiveness, we conducted experiments combining PersonaCraft with diverse style LoRAs, including Crayon [55], Pastel [8], 3D Render [27], Pixel Art [53], Illustration [7], Frosting Lane [5], Pokémon Trainer [67], JoJo [54], Graphic Novel [15], and Cartoon [6]. The results, shown in Fig. 8, highlight the method’s ability to adapt to different styles effectively. Notably, styles such as Pastel, Illustration, JoJo, and Pokémon Trainer introduce changes in facial and body characteristics, occasionally altering perceived identity, due to their bias. Nevertheless, the outcomes remain visually compelling and demonstrate the versatility of our approach.
4.5 Efficiency Analysis
As analyzed in Tab. 3, we compare the inference times for multi-identity personalized synthesis across different methods, specifically for generating images with three distinct identities. Textual Inversion [23] and DreamBooth [60], which rely on optimization-based personalization, require a batch size of 4 and 500 optimization steps per identity. This process results in significantly longer inference times, making these methods highly inefficient for real-time applications.
4.6 Ablation Studies
As shown in Tab. 4, the use of the body shape parameter enhances body shape preservation during personalization. Additionally, Fig. 9(a) demonstrates that our SCNet outperforms others in pose-guided image generation, effectively preserving human anatomy even in complex poses and occlusions. Moreover, Fig. 9(b) shows that our masking method improves performance, while adding features without masking increases feature variance and results in artifacts.
Additional results, including qualitative results, stylized artwork, PersonaCraft with LoRA and other face control modules, and ablation studies, can be found in the supplementary material.
5 Discussion

### v2

@@ 3.2 Occlusion-Focused Generation
$\footnotesize\mathbb{E}_{\bm{x}_{0},t,y,\bm{e}_{\text{SMPLx}},\bm{M}_{\text{occ}},\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y,\mathcal{E}^{\text{Occ}}_{\omega}\left(\bm{e}_{\text{SMPLx}}\odot\bm{M}_{\text{occ}}\right)\right)\right\|_{2}^{2}\right].$ (3)
Occlusion-Aware Classifier-Free Guidance. We observe that increasing classifier-free guidance (CFG) in occluded regions enhances anatomical consistency by strengthening 3D pose information, resolving local ambiguities. However, a uniformly high CFG strength induce over-saturation in non-occluded areas, as noted in prior works [68, 69, 47]. This highlights the need for a spatially adaptive guidance strategy [72] tailored for occlusion in human scene scenarios. To this end, we introduce Occlusion-Aware Classifier-Free Guidance (OccCFG) as shown in Fig. 3:
$\footnotesize\hat{\bm{\epsilon}}=\bm{\epsilon}_{\text{uncond}}+(k_{\text{occ}}\bm{M}_{\text{occ}}+k_{\text{base}}(1-\bm{M}_{\text{occ}}))(\bm{\epsilon}_{\text{cond}}-\bm{\epsilon}_{\text{uncond}})$ (4)
where $k_{\text{occ}}$ and $k_{\text{base}}$ are CFG scales for occluded and non-occluded regions, respectively, $\bm{M}_{\text{occ}}$ is the occlusion mask, $\bm{\epsilon}_{\text{uncond}}$ is the unconditional noise prediction, and $\bm{\epsilon}_{\text{cond}}$ is the conditional noise prediction.
3.3 Full-Body Personalized Generation
Another key advantage of SMPLx beyond its robust pose representation is its ability to encode body shape via shape coefficients, enabling a step forward in full-body multi-human personalization, in contrast to prior methods that focus solely on facial identity [43, 97, 31, 93, 85, 81, 90]. Furthermore, existing approaches struggle to disentangle garments from body shape, whereas ours explicitly isolates body shape as the true identity-defining feature.
Figure 7: Comparison of personalized multi-human scene generation with the methods which are not direct baselines. Yellow arrows highlight anatomical inconsistencies in poses and occlusions, while Pink arrows indicate identity mixing or duplication in the baselines.
Figure 8: Qualitative comparison of pose-controlled multi-human generation. yellow arrows highlight unnatural anatomical structures, with yellow boxes providing zoomed-in views.
Table 3: Quantitative evaluation of pose-controlled human generation across pose accuracy, text alignment, and image quality.
Pose-Controlled Human Generation Pose Text Image quality MPJPE (3D) ↓ AP-0.5 (2D) ↑ CLIP sim ↑ IS ↑ KID ↓ T2I Adpater-SDXL 198.14 0.323 0.269 3.740 0.129 ControlNet-SDXL 114.93 0.291 0.285 3.062 0.138 ControlNet-Flux 102.64 0.393 0.218 3.651 0.107 HumanSD-SD2 103.62 0.357 0.261 3.610 0.082 Ours 62.647 0.495 0.274 4.113 0.091
Table 4: Top-1 preference rates from the user study on naturalness, pose consistency, and text-image correspondence in pose-controlled multi-human scene generation.
Top 1 (%) Natural.↑ Pose consistency↑ Text corr.↑ T2I Adpater-SDXL 9.53 7.31 8.19 ControlNet-SDXL 17.72 18.19 19.36 Ours 72.75 74.50 72.46
The overall process of full-body personalized image synthesis is illustrated in Fig.5. Given full-body reference images $\{I_{\text{ref}}^{(i)}\}_{i=1}^{N_{\text{human}}}$ of $N_{\text{human}}$ individuals, target poses $\{\bm{p}^{(i)}\}_{i=1}^{N_{\text{human}}}$, and a text prompt $y$, our goal is to generate a personalized image $I_{P}$ that faithfully preserves both face and body identities while ensuring pose consistency. To achieve this, we use MultiHMR [12] to estimate SMPLx body shape parameters $\bm{\beta}^{(i)}$ and InsightFace [1] to extract face embeddings $\bm{f}^{(i)}$.
@@ 3.3 Full-Body Personalized Generation
We then leverage our depth-normal SCNet $\mathcal{E}^{\text{SC}}_{\phi_{D}}$-$\mathcal{E}^{\text{SC}}_{\phi_{N}}$, OccNet $\mathcal{E}^{\text{Occ}}_{\omega}$, and face ControlNet IdentityNet [82] $\mathcal{E}^{\text{ID}}_{\psi}$. Let $F^{(k)}_{\theta}(\cdot)$ denote the $k$-th neural block and $\bm{s}^{(k)}$ the corresponding input feature map. When $\bm{R}_{\text{sc}}=\mathcal{E}^{\text{SC}}_{\phi_{D}}(\bm{d}_{\text{SMPLx}})+\mathcal{E}^{\text{SC}}_{\phi_{N}}(\bm{n}_{\text{SMPLx}})$, $\bm{R}_{\text{occ}}=\mathcal{E}^{\text{Occ}}_{\omega}(\bm{d}_{\text{SMPLx}})$, and $\bm{R}^{(i)}_{\text{id}}=\mathcal{E}^{\text{ID}}_{\psi}(\bm{f}^{(i)},\bm{p}_{\text{face}}^{(i)})$ are the residual features from each model, the next feature map $\bm{s}^{(k+1)}$ is obtained by adding these residual features to the neural block output $F^{(k)}_{\theta}(\bm{s}^{(k)})$, scaled by their respective conditioning weights $\alpha_{\text{sc}}$, $\alpha_{\text{occ}}$, and $\alpha_{\text{id}}$, and modulated by the face masks $\bm{M}_{\text{face}}^{(i)}$ and occlusion mask $\bm{M}_{\text{occ}}$:
@@
$\footnotesize\begin{split}\bm{s}^{(k+1)}&=F^{(k)}_{\theta}(\bm{s}^{(k)})+\alpha_{\text{sc}}(1-\bm{M}_{\text{occ}})\bm{R}_{\text{sc}}\\ &\quad+\alpha_{\text{occ}}\bm{M}_{\text{occ}}\bm{R}_{\text{occ}}+\alpha_{\text{id}}\sum_{i=1}^{N_{\text{human}}}\bm{M}_{\text{face}}^{(i)}\ast\bm{R}^{(i)}_{\text{id}}.\end{split}$ (5)
Dual-Pathway Body Shape Personalization. We introduce a dual-pathway approach for body shape personalization, leveraging both SMPLx-based and text-based representations. Compared to textual descriptions of body shape, SMPLx-based representations provide more spatially fine-grained encoding of the identity. However, they are limited in accurately conveying body composition details, such as fat and muscle distribution [62]. In challenging cases where body shape characteristics conflict (e.g., a muscular physique with high body fat), our method optionally incorporates a textual pathway to enhance body shape representation. Specifically, a CLIP-based classifier [65] extracts body shape attributes in text, which are applied using regional prompting [10] and integrated with SCNet. Further details are provided in the supplementary material.
4 Experiments
Implementation Details. PersonaCraft is built on Stable Diffusion XL (SDXL)[64]. We extend the MPII dataset[7] with textual descriptions [38] and 3D human reconstructions from MultiHMR [12] using SMPLx [61]. SCNet and OccNet were fine-tuned on ControlNet [91] with a learning rate of 1e-5, Adam optimizer, batch size of 2, for 50,000 iterations on an NVIDIA A100 GPU. We use 30 denoising steps. We set the base CFG scale to 3 ($k_{\text{base}}=3$) and increased the scale for occluded regions to 5 ($k_{\text{occ}}=5$) to enhance occlusion handling. We set all conditioning scales ($\alpha_{\text{sc}}$, $\alpha_{\text{occ}}$, and $\alpha_{\text{id}}$) to 0.8 across all cases.
Figure 9: Effect of our occlusion-aware 3D pose & shape conditioning components. D and N denote the conditioning types of SCNet: depth, normal.
Figure 10: Effect of our occlusion-aware CFG (OccCFG). Each label refers to the CFG scale applied to specific regions. "Base=3" means the CFG scale is 3 for all regions, while "Base=3, Seg=5" indicates that the CFG scale is 5 for the human segmentation region. "Occ" refers to the occlusion mask region.
Table 5: Ablation study on SCNet, OccNet, and OccCFG for pose consistency. D, N, and R denote the conditioning types of SCNet: depth, normal, and RGB representations, respectively.
Ablation Study Pose MPJPE (3D)$\downarrow$ AP-0.5 (2D)$\uparrow$ SCNet-D 71.20 0.423 SCNet-N 79.30 0.398 SCNet-R 82.52 0.394 SCNet-D + SCNet-N 63.16 0.494 SCNet-D + SCNet-N + OccNet 62.92 0.499 SCNet-D + SCNet-N + OccNet + OccCFG 62.64 0.495
Metrics. For personalized multi-human scene generation, face identity preservation is measured using FaceNet [71], following FastComposer [85]. Body shape preservation is evaluated via cosine similarity of SMPLx body shape parameters $\bm{\beta}$. Pose consistency is assessed using MPJPE (3D)[35], while 2D pose estimation is evaluated with AP-0.5 (2D) by comparing the target pose to the estimated pose from generated images. Text-image alignment is measured via CLIP-L/14 similarity. Image quality is quantified using Inception Score (IS)[70] and Kernel Inception Distance (KID) [14]. A user study evaluates perceptual quality.
For pose-controlled multi-human scene generation, the same key metrics are applied: MPJPE (3D) and AP-0.5 (2D) for pose accuracy, CLIP similarity for text-image alignment, and IS/KID for image quality.
Test Dataset. We test on the COCO-WholeBody dataset [36], containing 1$\sim$5 people per image. A total of 1,000 images were selected (200 per group). Text prompts were extracted via BLIP [49]. MultiHMR [11] provided 2D poses and 3D conditioning.
Additional details on experiments, results (including qualitative results), stylized artwork, PersonaCraft with LoRA and other face control modules, as well as ablation studies, are provided in the supplementary material.
4.1 Personalized Multi-Human Scene Generation
Baselines. We mainly compared PersonaCraft with baseline methods for single-shot, multi-identity, and pose-controllable synthesis, all implemented using SDXL [64]. Key baselines include OMG [43] with InstantID/IPAdapter/IPAdapter-Face (IPA-Face) [81, 90], and 2D pose ControlNet [91]. We also evaluated optimization-based methods like DreamBooth [66] and Texture Inversion [22]. Both qualitative and quantitative comparisons, as well as a user study, were conducted. Additionally, qualitative comparisons were made with UniPortrait [31], MS-Diffusion [83], and FastComposer [86].
Qualitative Results. As shown in Fig. 6, PersonaCraft surpasses baselines in multi-identity scene generation, particularly in identity preservation and occlusion handling. Our 3D-aware conditioning and occlusion-focussed generation enhance the preservation of body shapes and identity features, and ensures pose fidelity, producing realistic generations with fewer artifacts. Baseline methods struggle with occlusion and body shape accuracy, often leading to distorted identities and unrealistic compositions, especially in multi-subject scenes. PersonaCraft excels in occlusion-aware full-body personalization, maintaining consistent identities across varying poses.
Quantitative Results. We evaluate PersonaCraft on face ID preservation, body shape, pose accuracy, text-image alignment, and image quality. As shown in Tab. 1, our method consistently outperforms baselines in identity and body shape preservation. It achieves the lowest MPJPE (3D) and highest AP (2D), indicating superior alignment with input poses. Furthermore, PersonaCraft outperforms baselines in IS and KID, producing perceptually aligned generations.
User Study. We conducted a user study evaluating naturalness, identity, body shape, and text-image correspondence across baselines and our method. With 18,540 responses from 103 participants, Tab. 2 shows our method achieved the highest Top-1 preference across all metrics, demonstrating superior perceptual quality in personalized multi-human scene generation.
4.2 Pose-Controlled Multi-Human Scene Generation
Baselines. Key baselines include ControlNet-SDXL [91, 64], T2I Adapter-SDXL [56], ControlNet-Flux [48, 91], and HumanSD [48]. Both qualitative and quantitative comparisons were conducted for all baselines. Additionally, a user study was performed specifically for SDXL-based baselines, including ControlNet-SDXL and T2I Adapter-SDXL.
Figure 11: Effectiveness of dual-pathway body shape personalization in challenging cases where body shape characteristics conflict (e.g., a muscular physique with high body fat).
Qualitative Results. We evaluate pose-controlled multi-human scene generation by comparing PersonaCraft with existing methods in both single-human and multi-human settings (Fig. 8). Our method achieves superior pose alignment while preserving identity consistency and body shape fidelity. Baselines relying on 2D pose conditioning (e.g., T2IAdapter-SDXL [56], ControlNet-SDXL [91, 64]) struggle with occlusion handling, leading to misaligned poses and distorted structures. ControlNet-Flux [48, 91] exhibits instability: at a lower conditioning scale (0.8), it fails to maintain anatomical coherence, while at a higher scale (1.0), it improves pose accuracy but reduces image fidelity. HumanSD [37] lacks precise pose control, producing unrealistic compositions. In contrast, PersonaCraft balances pose accuracy, identity preservation, and body realism, demonstrating the efficacy of our 3D-aware pose conditioning.
Quantitative Results. We evaluate pose-controlled multi-human scene generation in terms of pose accuracy, text-image correspondence, and image quality. As shown in Tab. 3, PersonaCraft achieves the lowest MPJPE (3D) and highest AP-0.5 (2D), demonstrating precise pose alignment and keypoint localization. Compared to ControlNet-based baselines, which suffer from structural inconsistencies and occlusion issues, our method maintains accurate pose fidelity while preserving identity and body shape. PersonaCraft also achieves the highest IS and lowest KID, indicating superior perceptual realism. While text-image correspondence scores are comparable across methods, prior studies [45, 42] highlight CLIP’s limitations in evaluating complex multi-human scenes. A user study further analyzes perceptual quality.
User Study. We conducted a user study to assess naturalness, pose consistency, and text-image correspondence. Participants ranked their top three preferences among T2I Adapter-SDXL, ControlNet-SDXL, and PersonaCraft, collecting 15,390 responses from 114 participants. As shown in Tab. 4, PersonaCraft achieved the highest Top-1 preference across all metrics, excelling in naturalness (72.75%), pose consistency (74.50%), and text alignment (72.46%). These results confirm that our occlusion-aware 3D pose conditioning significantly improves perceptual quality, producing well-posed, visually coherent multi-human scenes.
4.3 Ablation Studies and Analysis
Comparison of 3D Pose Representations for SCNet. In Tab. 5, we evaluate different SMPLx renderings—depth, normal, and RGB—as conditioning inputs for SCNet. We find that depth yields the best pose consistency among these representations. Additionally, combining the depth and normal representations achieves the best performance. Therefore, we adopt this combination as the base conditioning for SCNet.
Effectiveness of OccNet and OccCFG. In Tab. 5, we also evaluate the effectiveness of OccNet and OccCFG. We found that OccNet and OccCFG improves pose consistency. Also, as analyzed in Fig. 9, effect of our Occlusion-aware 3D pose & shape conditioning components. Using only 2D pose leads to front-back ambiguities and structural inconsistencies, while relying solely on 3D pose (SCNet) struggles with fine-grained occluded regions. Using only OccNet improves occluded region synthesis but fails to maintain overall pose alignment. Our full model (SCNet + OccNet + OccCFG) effectively preserves pose structure while handling occlusions. In Fig. 10, we present a detailed analysis of OccCFG. We observe that increasing classifier-free guidance (CFG) in occluded regions improves anatomical consistency by leveraging stronger 3D pose information, effectively resolving local ambiguities. However, uniformly high CFG strength leads to over-saturation in non-occluded areas. We found that applying CFG only in the human segmentation region also results in high CFG, whereas our OccCFG avoids issues in unoccluded regions while maintaining effective guidance in occluded areas.
Dual-Pathway Body Personalization. As shown in Fig 11, the SMPLx-based pathway captures fine-grained identity features but struggles with detailed body composition (e.g., muscle vs. fat distribution). The text-based pathway refines body shape using semantic attributes. Our approach combines both for more accurate and realistic personalization.
@@ 4.4 User-Defined Body Shape Control
A key feature of PersonaCraft is its user-defined body shape control, as shown in Fig. 12. Users can select a reference to adapt body characteristics for personalized images. The system also allows body proportions to be adjusted through interpolation or extrapolation between references, offering enhanced customization for precise body shape control.
5 Limitations